# A Study on the Calibration of In-context Learning

### 11 Dec 2023 v2

@@ Abstract
Modern auto-regressive language models are trained to minimize log loss on broad data by predicting the next token so they are expected to get calibrated answers in next-token prediction tasks. We study this for in-context learning (ICL), a widely used way to adapt frozen large language models (LLMs) via crafting prompts, and investigate the trade-offs between performance and calibration on a wide range of natural language understanding and reasoning tasks. We conduct extensive experiments to show that such trade-offs may get worse as we increase model size, incorporate more ICL examples, and fine-tune models using instruction, dialog, or reinforcement learning from human feedback (RLHF) on carefully curated datasets. Furthermore, we find that common recalibration techniques that are widely effective such as temperature scaling provide limited gains in calibration errors, suggesting that new methods may be required for settings where models are expected to be reliable.
@@ 1 Introduction
Language models (LMs) that encompass transformer-based architectures (Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023) can generate coherent and contextually relevant texts for various use cases. Despite their impressive performance, these models occasionally produce erroneous or overconfident outputs, leading to concerns about their calibration (Dawid, 1982; DeGroot and Fienberg, 1983) which measures how faithful a model’s prediction uncertainty is. Such a problem is pressing as users adapt them using a recent paradigm called in-context learning (Brown et al., 2020) to construct performant predictors, especially for applications in safety-critical domains (Bhatt et al., 2021; Kadavath et al., 2022; Pan et al., 2023).
We provide an in-depth evaluation and analysis of how well these models are calibrated - that is, the alignment between the model’s confidence in its predictions and the actual correctness of those predictions. This token-level calibration assessment will enable us to measure the discrepancy between the model’s perceived and actual performance through a Bayesian uncertainty lens, providing a valuable metric for assessing the model’s accuracy and reliability.
@@
(b) The accuracy and calibration of LLaMA-30B
Figure 1: The accuracy-calibration trade-off of in-context learning. (a) ICL concerns taking task-specific examples as the prompt to adapt a frozen LLM to predict the answer. (b) Classification accuracy and expected calibration error of ICL. As the number of ICL samples increases, the prediction accuracy improves (Left); at the same time, the calibration gets worse (Right).
We find that LMs including GPT-2 (Radford et al., 2019) and LLaMA (Touvron et al., 2023a) are poorly calibrated and there exists a calibration-accuracy trade-off (Fig.1), i.e. as we increase the amount of in-context samples, the prediction accuracy and calibration error both increase. Crucially, this calibration degradation worsens as the model size increases or when fine-tuning occurs using specialized data, such as curated instructions (Dubois et al., 2023), dialogues (Zheng et al., 2023), or human preference data (Ziegler et al., 2019). Though previous work (Braverman et al., 2020) shows the entropy of each generation step is drifting and can be recalibrated via scaling techniques (Platt et al., 1999) such as temperature scaling (Guo et al., 2017), we show that the miscalibration issue in ICL can not be easily addressed using such well-established recalibration approaches that rely on additional validation data.
Moreover, we study the trade-off in reasoning tasks that involve the generation of explanations (Camburu et al., 2018; Nye et al., 2021; Wei et al., 2022) before the answer, showing that the model can produce confidently wrong answers (using confidence histograms and reliability plots) when prompted with explanations on Strategy QA (Geva et al., 2021), Commonsense QA (Talmor et al., 2018), OpenBook QA (Mihaylov et al., 2018), World Tree (Jansen et al., 2018). We carefully design our human assessment to observe that, with the increase in model sizes and the quantity of ICL examples, there is a corresponding rise in the proportion of confidently predicted examples among those incorrectly forecasted. Moreover, we find that a high proportion of wrong predictions are of high confidence and showcase those typical confidently wrong examples of LLMs.
In-context learning has been hypothesized as meta-learning by gradient descent in their forward pass (Von Oswald et al., 2023), which might hopefully yield calibrated predictions (Błasiok et al., 2023) if the models are getting close to local optimality with respect to test loss through meta-optimization. However, the fact that choosing ICL samples from the validation set does not naturally lead to calibrated predictions shows that ICL learns in a fairly different way than SGD. We design controlled experiments to illustrate task learning properties of ICL, showing that when examples in the prompt demonstrate consistent task properties, the learning performance would be improved while the accuracy-calibration trade-off still persists.
@@ 2 Related Work
Calibration of LLMs. Calibration is a safety property to measure the faithfulness of machine learning models’ uncertainty, especially for error-prone tasks using LLMs. Previous works find that pre-training (Desai and Durrett, 2020) and explanation (Zhang et al., 2020; González et al., 2021) improves calibration. Models can be very poorly calibrated when we prompt LMs (Jiang et al., 2021), while calibration can also depend on model size (Kadavath et al., 2022). (Braverman et al., 2020) assesses the long-term dependencies in a language model’s generations compared to those of the underlying language and finds that entropy drifts as models such as GPT-2 generate text. The intricacy of explanations on complementary team performance poses additional challenges due to the overreliance on explanations of users regardless of their correctness (Bansal et al., 2021). (Mielke et al., 2022) gives a framework for linguistic calibration, a concept that emphasizes the alignment of a model’s expressed confidence or doubt with the actual accuracy of its responses. The process involves annotating generations with <DK>, <LO>, <HI> for confidence levels, then training the confidence-controlled model by appending the control token <DK/LO/HI> at the start of the output, followed by training a calibrator to predict these confidence levels, and finally predicting confidence when generating new examples. (Tian et al., 2023) finds that asking LLMs for their probabilities can be better than using conditional probabilities in a traditional way. (Shih et al., 2023) proposes a simple amortized inference trick for temperature-scaled sampling from LMs and diffusion models. To enhance the estimation of uncertainty in language models, (Kuhn et al., 2023) developed a method that aggregates log probabilities across semantically equivalent outputs. This approach utilizes bidirectional entailment through a model to identify outputs that are semantically similar, thereby refining the uncertainty estimation process. (Cole et al., 2023) identifies the calibration challenge in ambiguous QA and distinguishes uncertainty about the answer (epistemic uncertainty) from uncertainty about the meaning of the question (denotational uncertainty), proposing sampling and self-verification methods. (Kamath et al., 2020) trains a calibrator to identify inputs on which the QA model errs and abstains when it predicts an error is likely. (Zhao et al., 2023) proposes the Pareto optimal learning assessed risk score for calibration and error correction but requires additional training. (Kalai and Vempala, 2023) shows the trade-off between calibration and hallucination but they didn’t study it in an ICL setting and how the predicted answer’s accuracy would impact those two safety aspects.
@@
In-context Learning. Large models such as GPT-3 (Brown et al., 2020) have demonstrated the potential of in-context learning, a method where the model infers the task at hand from the context provided in the input, without requiring explicit retraining or fine-tuning for each new task. Some recent works attempt to understand ICL through meta-learning (Von Oswald et al., 2023), Bayesian inference (Xie et al., 2021), mechanistic interpretability (Olsson et al., 2022), algorithm selection (Bai et al., 2023), synthetic data and simple function classes (Garg et al., 2022; Akyürek et al., 2022). Notably, unlike previous works (Zhao et al., 2021; Han et al., 2023; Fei et al., 2023; Zhou et al., 2023a) that focus on improving task accuracy using the same “calibration” terminology, we study the uncertainty of ICL and measure its trade-off with accuracy.
@@ 3 Background
Table 1: Accuracy and Calibration of LLaMA-30B model with three sizes across four text classification datasets and four reasoning datasets
@@
0-shot 1-shot 4-shot 8-shot
Acc ECE Acc ECE Acc ECE Acc ECE
@@
Strategy QA 0.452 0.039 0.617 0.047 0.679 0.045 0.678 0.088
Commonsense QA 0.354 0.143 0.613 0.268 0.642 0.279 0.733 0.313
World Tree 0.53 0.253 0.594 0.276 0.655 0.31 0.24 0.230
OpenBook QA 0.388 0.122 0.55 0.15 0.641 0.226 0.66 0.246
@@
We extract token-level answer probabilities of LLMs, [^1] We also normalize the probability $\mathcal{P}_{\theta}\left(y_{n}\mid w_{<n}\right)\in\Delta^{K}$ for classification problems with $K$ choices e.g. for binary classification tasks, we filter and extract probabilities $P(``\textrm{Yes}")$ and $P(``\textrm{No}")$, based on which we calculate the following statistics for studying the confidence and calibration of LMs:
@@
Entropy rate. We denote the entropy of a token $w_{t}$ at position $t$ as $H(w_{t}|w_{<t})=-\mathbb{E}_{w_{t}\sim\mathcal{P}_{\theta}(\cdot|w_{<t})}[\log\mathcal{P}_{\theta}(w_{t}|w_{<t})]$. We typically measure it based on the answer token via setting $w_{t}=y_{n}$. Note that auto-regressive LLMs are trained via maximizing the negative log-likelihood objective $\mathcal{L}=-\mathbb{E}_{t}[\log\mathcal{P}_{\theta}\left(w_{t}|w_{<t}\right)]$ on massive corpora.
@@
4 Experimental Results
@@ 4.1 Experimental Settings
Models. We study decoder-only autoregressive LMs involving LLaMA (Touvron et al., 2023a), and its variants fine-tuned with instruction, dialog, or RLHF like Alpaca (Dubois et al., 2023), Vicuna (Zheng et al., 2023), and LLaMA2-Chat (Touvron et al., 2023b).
@@ 4.2 Numerical Results
The performance of LLaMA. We record the performance and calibration errors for $k$-shot learning ($k=0,1,4,8$), characterizing the calibration-accuracy trade-off in both classic and realistic settings (Tab. 1). Our findings are two-fold: as more in-context examples are included, we observe a concurrent rise in both accuracy and calibration error across most situations, indicating a diminishing reliability of token-level probability in high performance ICL regimes; moreover, when $k=0$ increases to $k=1$, there is a marked boost in both accuracy and calibration error, demonstrating the importance of in-context examples and the significance of the established trade-off. In particular, for reasoning tasks, we explore prompting approaches that explicitly include explanations in reasoning tasks, i.e. scratchpad (Nye et al., 2021) or chain-of-thought (Wei et al., 2022), showing that calibration degrades after generating a long context for reasoning and explaining the final answer.
Table 2: Accuracy and calibration of LLaMA-7B model with different prompt repetition strategies.
Model Strategy 1-shot 4-shot 8-shot AVG ACC AVG ECE
Acc ECE Acc ECE Acc ECE
LLaMA-7B Repeat-context 78.200 0.281 78.370 0.283 77.470 0.274 78.013 0.279
Repeat-prompt 78.200 0.281 89.470 0.395 90.330 0.403 86.000 0.360
Normal 78.200 0.281 95.200 0.452 96.130 0.462 89.843 0.398
The effect of temperature scaling. We experiment with three strategies in applying temperature scaling methods (Guo et al., 2017) to fix miscalibration:
@@
    We learn one temperature for each $k$-shot ICL, i.e., we learn different temperatures for different shot numbers in ICL;
@@
    Learn a temperature from the training split (zero-shot) and apply it to all test samples with different shot numbers;
@@
    For each experiment, we fix the prompt and learn the temperature for the fixed prompt. That is, for every possible ICL prompt, we learn a corresponding temperature for calibration.
@@
Figure 3: Accuracy and calibration errors of LLaMA and its fine-tuned variants.
Looking into Fig. 6, none of the above strategies achieves satisfactory calibration performance, which is in contrast to the well-studied supervised learning setting where scaling the confidence scores (via temperature scaling) can effectively reduce calibration errors (Guo et al., 2017). The fact that applying a post-processing calibration method, such as temperature scaling, cannot directly resolve the miscalibration issue suggests that ICL might have substantially different properties compared to making predictions via classical supervised learning models.
@@
The effect of fine-tuning. We show that vicuna, alpaca and LLaMA2-Chat are all more accurate but less calibrated than their LLaMA counterpart backbones (Fig. 3), the margin is especially large for reasoning tasks and vicuna. Our finding indicates that fine-tuning might significantly degrade calibration, corroborating the evidence reported in GPT-4 (OpenAI, 2023), albeit it can improve the reasoning accuracy dramatically. Our results provide evidence that though fine-tuned on carefully curated datasets can greatly improve question-answering performance, especially for hard tasks like reasoning problems, attention may need to be paid when assessing the calibration of those models’ predictions. We also note that the pattern of zero-shot performance is fairly different for two fine-tuned models, i.e. vicuna, and alpaca.
The effect of prompt formats. In our study, we explore the effects of different prompt strategies using three distinct methods. First, the Repeat-context approach involves constructing prompts as $w_{0}=[x_{1},x_{1},...,x_{1},y_{1}]$, where the context $x_{1}$ is repeated n times, but the label $y_{1}$ is not included in the repetition. Next, the Repeat-prompt strategy shapes the prompt as $w_{0}=[x_{1},y_{1},...,x_{1},y_{1}]$, where both the context $x_{1}$ and the label $y_{1}$ are repeated n times. Finally, the Normal involves constructing the prompt as $w_{0}=[x_{1},y_{1},x_{2},y_{2},...,x_{n-1},y_{n-1},x_{n},y_{n}]$, systematically incorporating distinct context-label pairs.
The findings, as detailed in Tab. 2, reveal certain insights: firstly, integrating labels within prompts significantly decreases uncertainty and enhances reasoning performance. In contrast, simply repeating the context without labels does not lead to better outcomes. Secondly, the diversity of ICL examples in prompt construction notably affects performance, especially for larger language models. Those observations corroborate that ICL is performant when the number of ICL examples is large and they demonstrate consistent task properties. Importantly, the trade-off persists for such controlled scenarios, i.e. improving learning performance in either way would suffer from poorer calibration.

### 02 Jan 2024 v3

@@ Abstract
Accurate uncertainty quantification is crucial for the safe deployment of language models (LMs), and prior research has demonstrated improvements in the calibration of modern LMs. Our study focuses on in-context learning (ICL), a prevalent method for adapting static LMs through tailored prompts, and examines the balance between performance and calibration across a broad spectrum of natural language understanding and reasoning tasks. Through comprehensive experiments, we observe that, with an increasing number of ICL examples, models initially exhibit increased miscalibration before achieving better calibration and miscalibration tends to arise in low-shot settings. Moreover, we find that methods aimed at improving usability, such as fine-tuning and chain-of-thought (CoT) prompting, can lead to miscalibration and unreliable natural language explanations, suggesting that new methods may be required for scenarios where models are expected to be reliable.
@@ 1 Introduction
Language models (LMs) that encompass transformer-based architectures (Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023) can generate coherent and contextually relevant texts for various use cases. Despite their impressive performance, these models occasionally produce erroneous or overconfident outputs, leading to concerns about their calibration (Dawid, 1982; DeGroot and Fienberg, 1983) which measures how faithful a model’s prediction uncertainty is. Such a problem is pressing when users adapt them using a recent paradigm called in-context learning (Brown et al., 2020) to construct performant predictors, especially for applications in safety-critical domains (Bhatt et al., 2021; Kadavath et al., 2022; Pan et al., 2023).
We provide an in-depth evaluation and analysis of how well these models are calibrated - that is, the alignment between the model’s confidence in its predictions and the actual correctness of those predictions. This token-level calibration assessment enables us to measure the discrepancy between the model’s perceived and actual performance through a Bayesian uncertainty lens, providing a valuable metric for assessing the model’s accuracy and reliability.
@@
(b) The accuracy and calibration of LLaMA-7B
Figure 1: The accuracy-calibration trade-off of in-context learning. (a) ICL concerns taking task-specific examples as the prompt to adapt a frozen LLM to predict the answer. (b) Classification accuracy and expected calibration error of ICL. As the number of ICL samples increases, the prediction accuracy improves (Left); at the same time, the calibration first worsens ($k<3$) and then becomes better (Right).
We find that LM such as LLaMA (Touvron et al., 2023a) is poorly calibrated in performant settings and there exists a calibration-accuracy trade-off (Fig.1) for low-shot settings ($k<4$), i.e. as we increase the amount of in-context samples, both prediction accuracy and calibration error increase. Such a trade-off can be improved using more ICL examples ($k=8$) and larger models. Crucially, this calibration degradation worsens when fine-tuning occurs using specialized data to improve usability, such as curated instructions (Dubois et al., 2023), dialogues (Zheng et al., 2023), or human preference data (Ziegler et al., 2019). Though previous work (Braverman et al., 2020) shows the entropy of each generation step is drifting and can be recalibrated via scaling techniques (Platt et al., 1999) such as temperature scaling (Guo et al., 2017), we show that in contrast to classic regimes, the miscalibration issue in ICL can not be easily addressed using such well-established recalibration approaches that rely on additional validation data.
Moreover, we study the trade-off in reasoning tasks that involve generation of explanations (Camburu et al., 2018; Nye et al., 2021; Wei et al., 2022) before the answer, showing that the model can produce confidently wrong answers (using confidence histograms and reliability plots) when prompted with explanations on Strategy QA (Geva et al., 2021), Commonsense QA (Talmor et al., 2018), OpenBook QA (Mihaylov et al., 2018), World Tree (Jansen et al., 2018). We carefully design our human evaluation and observe that, with the increase in model sizes and the quantity of ICL examples, there is a corresponding rise in the proportion of confidently predicted examples among those incorrectly forecasted. Moreover, we find that a high proportion of wrong predictions are of high confidence and showcase those typical confidently wrong examples of LMs.
In-context learning has been hypothesized as meta-learning by gradient descent in their forward pass (Von Oswald et al., 2023), which might hopefully yield calibrated predictions (Błasiok et al., 2023) if the models are getting close to local optimality with respect to test loss through meta-optimization. However, the fact that choosing ICL samples from the validation set does not naturally lead to calibrated predictions shows that ICL learns in a fairly different way than SGD. We design controlled experiments to illustrate the task learning properties of ICL, showing that when examples in the prompt demonstrate consistent task properties, the learning performance would be improved.
@@ 2 Related Work
Calibration of language models. Calibration is a safety property to measure the faithfulness of machine learning models’ uncertainty, especially for error-prone tasks using LLMs. Previous works find that pre-training (Desai and Durrett, 2020) and explanation (Zhang et al., 2020; González et al., 2021) improves calibration. Models can be very poorly calibrated when we prompt LMs (Jiang et al., 2021), while calibration can also depend on model size (Kadavath et al., 2022). (Braverman et al., 2020) assesses the long-term dependencies in a language model’s generations compared to those of the underlying language and finds that entropy drifts as models such as when GPT-2 generates text. The intricacy of explanations on complementary team performance poses additional challenges due to the overreliance on explanations of users regardless of their correctness (Bansal et al., 2021). (Mielke et al., 2022) gives a framework for linguistic calibration, a concept that emphasizes the alignment of a model’s expressed confidence or doubt with the actual accuracy of its responses. The process involves annotating generations with <DK>, <LO>, <HI> for confidence levels, then training the confidence-controlled model by appending the control token <DK/LO/HI> at the start of the output, followed by training a calibrator to predict these confidence levels, and finally predicting confidence when generating new examples. (Tian et al., 2023) finds that asking LMs for their probabilities can be better than using conditional probabilities in a traditional way. LHTS (Shih et al., 2023) is a simple amortized inference trick for temperature-scaled sampling from LMs and diffusion models. To aggregate log probabilities across semantically equivalent outputs, Kuhn et al. (2023) utilize bidirectional entailment through a model to identify outputs that are semantically similar, thereby refining the uncertainty estimation process. (Cole et al., 2023) identifies the calibration challenge in ambiguous QA and distinguishes uncertainty about the answer (epistemic uncertainty) from uncertainty about the meaning of the question (denotational uncertainty), proposing sampling and self-verification methods. (Kamath et al., 2020) trains a calibrator to identify inputs on which the QA model errs and abstains when it predicts an error is likely. (Zhao et al., 2023) proposes the Pareto optimal learning assessed risk score for calibration and error correction but requires additional training. (Kalai and Vempala, 2023) show the trade-off between calibration and hallucination but they didn’t study it in a realistic setting and how the predicted answer’s accuracy would impact those two safety aspects.
@@
In-context Learning. Large models such as GPT-3 (Brown et al., 2020) have demonstrated the potential of in-context learning, a method where the model infers the task at hand from the context provided in the input, without requiring explicit retraining or fine-tuning for each new task. Some recent works attempt to understand ICL through meta-learning (Von Oswald et al., 2023), Bayesian inference (Xie et al., 2021), mechanistic interpretability (Olsson et al., 2022), algorithm selection (Bai et al., 2023), synthetic data and simple function classes (Garg et al., 2022; Akyürek et al., 2022; Raventós et al., 2023). Notably, unlike previous works (Zhao et al., 2021; Han et al., 2023; Fei et al., 2023; Zhou et al., 2023a) that focus on improving task accuracy using the same “calibration” terminology, we study the uncertainty of ICL and measure its trade-off with accuracy.
@@ 3 Background
Table 1: Accuracy and Calibration of LLaMA-30B model with three sizes across four text classification datasets and four reasoning datasets. Results are excluded when the data exceeds the context length limit.
@@
0-shot 1-shot 2-shot 3-shot 4-shot 8-shot
ECE Acc ECE Acc ECE Acc ECE Acc ECE Acc ECE Acc
@@
Strategy QA 0.204 0.450 0.154 0.619 0.174 0.654 0.172 0.660 0.161 0.672 0.152 0.665
Commonsense QA 0.048 0.356 0.232 0.589 0.290 0.608 0.253 0.675 0.283 0.644 0.289 0.653
World Tree 0.112 0.534 0.211 0.570 0.251 0.621 0.185 0.680 0.206 0.646 - -
OpenBook QA 0.036 0.386 0.231 0.561 0.255 0.604 0.207 0.644 0.206 0.648 0.191 0.662
@@
We extract answer token probabilities of LMs, e.g. for binary classification tasks, we filter and extract probabilities $P(``\textrm{Yes}")$ and $P(``\textrm{No}")$, based on which we calculate the following statistics for studying the confidence and calibration of LMs:
@@
Entropy rate. We denote the entropy of a token $w_{t}$ at position $t$ as $H(w_{t}|w_{<t})=-\mathbb{E}_{w_{t}\sim\mathcal{P}_{\theta}(\cdot|w_{<t})}[\log\mathcal{P}_{\theta}(w_{t}|w_{<t})]$. We typically measure it based on the answer token via setting $w_{t}=y_{n}$. Note that auto-regressive LMs are trained via maximizing the negative log-likelihood objective $\mathcal{L}=-\mathbb{E}_{t}[\log\mathcal{P}_{\theta}\left(w_{t}|w_{<t}\right)]$ on massive corpora.
@@
4 Experiments
We briefly summarize our results and findings before explaining the experimental settings.
  • For the base LMs we considered, they are calibrated when prompting with a sufficient amount of ICL examples to get non-trivial performance.
  • As we increase the number of ICL examples, models tend to be first more miscalibrated and then calibrated. In low-shot settings ($k<4$), models can be mis-calibrated, in part due to poor data (aleatoric) uncertainty.
  • Interventions that improve usability such as fine-tuning, and chain-of-thought (CoT) prompting would lead to miscalibration. The generated explanations from CoT can improve predictive results but may not be reliable by human evaluation.
@@ 4.1 Experimental Settings
Models. We study decoder-only autoregressive LMs involving LLaMA (Touvron et al., 2023a), ranging from 7B to 30B, and its variants fine-tuned with instruction, dialog, or RLHF like Alpaca (Dubois et al., 2023), Vicuna (Zheng et al., 2023), and LLaMA2-Chat (Touvron et al., 2023b).
@@ 4.2 Numerical Results
Performance of LLaMA. We record the performance and calibration errors for $k$-shot learning ($k=0,1,2,3,4,8$), characterizing the calibration-accuracy trade-off in both classic and realistic settings (Tab. 1). Our findings are two-fold: as more in-context examples are included, we observe a concurrent rise in both accuracy and calibration error across most low-shot situations. Especially, when $k=0$ increases to $k=1$, there is a marked boost in both accuracy and calibration error, demonstrating the importance of in-context examples in learning performance while one single example may not be able to reduce aleatoric uncertainty. Moreover,
In particular, for reasoning tasks, we explore prompting approaches that explicitly include explanations in reasoning tasks, i.e. scratchpad (Nye et al., 2021) or chain-of-thought (Wei et al., 2022), showing that calibration significantly degrades after generating a long context for reasoning and explaining the final answer.
Table 2: Acc and ECE of LLaMA-7B model on SST-2 with different prompt repetition strategies.
1-shot 2-shot 4-shot 8-shot Avg Acc Avg ECE
ACC ECE ACC ECE ACC ECE ACC ECE
Vanilla 0.740 0.098 0.877 0.132 0.917 0.108 0.954 0.064 0.872 0.100
Repeat prompt 0.740 0.098 0.693 0.155 0.801 0.117 0.820 0.111 0.764 0.120
Repeat context 0.740 0.098 0.668 0.208 0.657 0.220 0.607 0.219 0.668 0.186
Post-hoc recalibraiton. We experiment with three strategies (Algorithm 1) in applying temperature scaling methods (Guo et al., 2017) to fix miscalibration:
@@
    (00-shot) Learn a temperature from the training split and apply it to all test samples with different shot numbers;
@@
    ($k$-shot) We learn one temperature for each $k$-shot ICL, i.e., we learn different temperatures for different shot numbers in ICL;
@@
    (Fix $w$) For each experiment, we fix the prompt and learn the temperature for the fixed prompt. That is, for every possible ICL prompt, we learn a corresponding temperature for calibration.
The training algorithm is shown in Appendix Alg. 1. Looking into Tab. 3, none of the above strategies achieves satisfactory calibration performance, which is in contrast to the well-studied supervised learning setting where scaling the confidence scores can effectively reduce calibration errors (Guo et al., 2017). The fact that applying a post-processing calibration method, such as temperature scaling, cannot directly resolve the miscalibration issue suggests that ICL might have substantially different properties compared to making predictions via classical supervised learning models.
@@
Figure 3: Accuracy and calibration errors of base models LLaMA and Mistral, as well as fine-tuned variants. The reported values for Acc and ECE represent averages across experiments conducted with 0, 1, …, 8 shots.
Table 3: ECE for different calibration strategies of base models LLaMA-7B across various shot settings.
Dataset Strategy Shot Avg
0 1 2 3 4 8
SST-2 Vanilla 0.038 0.098 0.133 0.136 0.124 0.064 0.099
00-shot 0.026 0.157 0.265 0.275 0.282 0.222 0.205
$k$-shot 0.026 0.104 0.103 0.101 0.087 0.035 0.076
Fix $w$ 0.027 0.079 0.056 0.051 0.051 0.030 0.049
CB Vanilla 0.133 0.274 0.164 0.148 0.192 0.206 0.186
00-shot 0.126 0.272 0.143 0.142 0.183 0.177 0.174
$k$-shot 0.126 0.247 0.137 0.153 0.203 0.251 0.186
Fix $w$ 0.138 0.183 0.127 0.141 0.150 0.194 0.155
RTE Vanilla 0.130 0.103 0.122 0.147 0.130 0.083 0.119
00-shot 0.117 0.100 0.130 0.154 0.130 0.094 0.121
$k$-shot 0.114 0.106 0.130 0.161 0.121 0.104 0.123
Fix $w$ 0.114 0.078 0.097 0.104 0.079 0.079 0.092
AGNews Vanilla 0.086 0.077 0.166 0.203 0.088 0.063 0.114
00-shot 0.101 0.087 0.149 0.198 0.117 0.078 0.122
$k$-shot 0.114 0.085 0.161 0.203 0.075 0.067 0.117
Fix $w$ 0.132 0.081 0.113 0.126 0.089 0.079 0.103
@@
The effect of fine-tuning. We show that vicuna, alpaca, and LLaMA2-Chat are all more accurate but less calibrated than their LLaMA counterpart backbones (Fig. 3), the margin is especially large for reasoning tasks and vicuna. Our finding indicates that fine-tuning might significantly degrade calibration, corroborating the evidence reported in GPT-4 (OpenAI, 2023), albeit it can improve the reasoning accuracy dramatically. Our results provide evidence that though fine-tuned on carefully curated datasets can greatly improve question-answering performance, especially for hard tasks like reasoning problems, attention may need to be paid when assessing the calibration of those models’ predictions. We also note that the pattern of zero-shot performance is fairly different for two fine-tuned models, i.e. vicuna, and alpaca. Moreover, we include results of Mistral-7B (Jiang et al., 2023), a sparse Mixture of Experts (MoEs) architecture with sliding window attention. As a base model, it shows similar performance and calibration compared with LLaMA2-7B, indicating that our conclusion still holds for the model pre-trained with significantly different data and architectures.
The effect of prompt formats. In our study, we explore the effects of different prompt strategies using three distinct methods. We consider predicting the label $y_{n}$ of test example $x_{n}$. First, the Repeat-context approach involves constructing prompts as $w_{0}=[x_{1},x_{1},...,x_{1},y_{1},x_{n}]$, where the context $x_{1}$ is repeated n-1 times, but the label $y_{1}$ is not included in the repetition. Next, the Repeat-prompt strategy shapes the prompt as $w_{0}=[x_{1},y_{1},...,x_{1},y_{1},x_{n}]$, where both the context $x_{1}$ and the label $y_{1}$ are repeated n-1 times. Finally, the Normal involves constructing the prompt as $w_{0}=[x_{1},y_{1},x_{2},y_{2},...,x_{n-1},y_{n-1},x_{n}]$, systematically incorporating distinct context-label pairs.
The findings, as detailed in Tab. 2, reveal certain insights: firstly, integrating labels within prompts significantly decreases uncertainty and enhances learning performance. In contrast, simply repeating the context without labels does not lead to better outcomes. Secondly, the diversity of ICL examples in the prompt greatly affects performance, a potential explanation is it promotes better task learning (Pan, 2023). Those observations corroborate that ICL is performant when the number of ICL examples is large and they demonstrate consistent task properties. Importantly, the trade-off persists for different controlled scenarios, i.e. as we increase the number of ICL examples, models tend to be first more miscalibrated and then calibrated.